Sorry about the earlier typo: The question was written as f(x)= x^2-10x+21
Answer by RAY100(1637)

f(x) =y = x^2 -10x +21
.
y= (x-3)(x-7)
.
@ y=0,,,,,,,,x-3=0,,x=3,,,,,,x-7=0,,,x=7